# Risk-Averse Total-Reward Reinforcement Learning

### v1

@@ 3.2 EVaR Q-learning Algorithm
In this section, we adapt our ERM Q-learning algorithm to compute the optimal policy for the static EVaR-TRC objective in (3). As mentioned in Section 2, EVaR is preferable to ERM because it is coherent and closely approximates VaR and CVaR. The construction of an optimal EVaR policy follows the standard methodology proposed in prior work su2025risk ; hau2023entropic . We only briefly summarize it in this paper.
@@
$h(\pi,\beta)$ $\;:=\;\operatorname{ERM}_{\beta}^{\pi,\mu}\left[q^{\pi}(\tilde{s}_{0},\tilde{a}_{0},\beta)\right]+\beta^{-1}\log(\alpha)$ (12)
@@ 1 Construct $\mathcal{B}(\beta_{0},\delta)$ as described in (13) ;
2 Compute $(\pi^{\star}_{\beta},\,h^{\star}(\beta))$ by Algorithm 1 for each $\beta\in\mathcal{B}(\beta_{0},\delta)$ where $h^{\star}(\beta)=\max_{\pi\in\Pi}h(\pi,\beta)$;
@@
Algorithm 2 summarizes the procedure for computing a $\delta$-optimal EVaR policy. Note that the value of $h(\pi,\beta)$ may be $\infty$- ∞, indicating either the ERM-TRC objective is unbounded for $\beta$, or the Q-learning algorithm diverged by random chance.
@@
Note that Algorithm 2 assumes a small $\beta_{0}$ as its input. The following remark gives the procedure we use to compute the value $\beta_{0}$. It is unclear whether obtaining an apriori bound on $\beta_{0}$ without knowing the model is possible. We, therefore, employ the heuristic outlined in Algorithm 3 that estimates a lower bound $x_{\min}$ and an upper bound $x_{\max}$ on the random variable of returns. Then, we set $\beta_{0}:=\nicefrac{{8\delta}}{{(x_{\max}-x_{\min})^{2}}}.$
@@ 4 Convergence Analysis
This section presents our main convergence guarantees for the ERM-TRC Q-learning and EVaR-TRC Q-learning algorithms. The proofs for this section are deferred to Appendix B.
We require the following standard assumption to prove the convergence of the proposed Q-learning algorithms. The intuition of 4.1 is that each state-action pair must be visited infinitely often.
@@
The input to Algorithm 1 and Algorithm 2 satisfies that
@@
The following theorem shows that the proposed ERM-TRC Q-learning algorithm enjoys convergence guarantees comparable to standard Q-learning. As Remark 2 states, $z_{\min}$ and $z_{\max}$ are chosen to ensure that $q$ values are bounded.
@@
For $\beta\in\mathcal{B}$, assume that the sequence $\left(\tilde{\eta}_{i}\right)_{i=0}^{\infty}$ and $\left((\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})\right)_{i=0}^{\infty}$ used in Algorithm 1 satisfies 4.1 and step size condition
@@
where $i\in\{i\in\mathbb{N}\mid(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\}$, if $\tilde{z}_{i}\in[z_{\min},z_{\max}]$ almost surely, then the sequence $(\tilde{q}_{i})_{i=0}^{\infty}$ produced by Algorithm 1 convergences almost surely to $q_{\infty}$ such that $q_{\infty}=\hat{B}_{\beta}q_{\infty}$.
The proof of Theorem 4.2 follows an approach similar to that in the proofs of standard Q-learning bertsekas1996neuro with three main differences. First, the algorithm converges for an undiscounted MDP, and $\hat{B}$ is not a contraction captured by (bertsekas1996neuro, , assumption 4.4), which is restated here as B.2. Instead, we show that $\hat{B}$ satisfies monotonicity in Lemma 4.3. Second, B.2 leads to a convergence result somewhat weaker than the results for a contraction. Then, a separate boundedness condition(bertsekas1996neuro, , proposition 4.6) is imposed. Third, using exponential loss function $\ell_{\beta}$ requires a more careful choice of step-size than the standard analysis.
@@
The following lemma shows that the loss function $l_{\beta}$ is strongly convex and has a Lipschitz-continuous gradient. These properties are instrumental in showing the uniqueness of our value functions and analyzing the convergence of our Q-learning algorithms by analyzing them as a form of stochastic gradient descent.
@@
Corollary 4.5, which follows immediately from Theorem 3.4, shows that Algorithm 2 enjoys convergence guarantees and converges to a $\delta$-optimal EVaR policy.
@@
For $\alpha\in(0,1)$ and $\delta>0$, assume that the sequence $\{\tilde{\eta}_{i}\}_{i=0}^{\infty}$ and $\{(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})\}_{i=0}^{\infty}$ used in Algorithm 1 satisfies 4.1 and step size condition
@@
where $i\in\{i\in\mathbb{N}\mid(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\}$, if $\tilde{z}_{i}\in[z_{\min},z_{\max}]$. Then Algorithm 2 converges to the $\delta$-optimal stationary policy $\pi^{\star}$ for the EVaR-TRC objective in (3) almost surely.
The proof of Corollary 4.5 follows a similar approach to the proofs of Theorem 4.2. However, we also need to show that one can obtain an optimal ERM policy for an appropriately chosen $\beta$ that approximates an optimal EVaR policy arbitrarily closely.
@@ 5 Numerical Evaluation
In this section, we evaluate our algorithms on two tabular domains: cliff walking (CW) andrew2018reinforcement and gambler’s ruin (GR) su2025risk .
First, to evaluate our EVaR Q-learning algorithm with risk levels $\alpha=0.2$ and $\alpha=0.6$, we use the cliff walk problem. In this problem, an agent starts with a random state (cell in the grid world) that is uniformly distributed over all non-sink states and walks toward the goal state labeled by $g$ shown in Figure 2. At each step, the agent takes one of four actions: up, down, left, or right. The action will be performed with a probability of $0.91$, and the other three actions will also be performed separately with a probability of $0.03$. If the agent hits the wall, it will stay in its place. The reward is zero in all transitions except for the states marked with $c$, $d$, and $g$. The agent receives $2$ in state $g$, $0.004$ in state $d$, and $-0.5,-0.6,-0.7,-0.8$, or $-0.9$ in cliff region marked with $c$ from the left to the right. If an agent steps into the cliff region, it will immediately be returned to the state marked with $b$. Note that the agent has unlimited steps, and the future rewards are not discounted.
@@
Before computing the optimal EVaR policy, we set $\beta_{c}=1^{-10}$ in Algorithm 3 to compute $c$ and $d$ in Remark 2. $c$ and $d$ are used to estimate the bounds of $\tilde{z}_{i}(\beta)$ in ERM-TRC Q-learning algorithm. Figure 2 shows the optimal policy for EVaR risk level $\alpha=0.2$ computed by Algorithm 2. Each cell represents a state, and the arrow direction indicates the optimal action to take in that state. Note that the optimal actions for states $c$ and $g$ are the same and are omitted in Figure 2 and Figure 2. Since the optimal policy is stationary, it can be interpreted and analyzed visually. The agent moves right to the last column of the grid world and then moves down to reach the goal. When the agent is close to the cliff region, it moves up to avoid risk. Figure 2 shows the optimal policy for EVaR risk level $\alpha=0.6$ computed by Algorithm 2. Different optimal actions are highlighted in blue. As we can see, for different risk levels $\alpha$, when the agent is near the cliff region, it exhibits consistent behaviors to avoid falling off the cliff. For the remaining states, the agent behaves differently.
@@
Figure 4: Mean and standard deviation of EVaR values with $\alpha=0.3$ on CW domain
Second, to understand the impact of risk aversion on the structure of returns, we simulate the two optimal EVaR policies over $48,000$ episodes and display the distribution of returns in Figure 4. The $x$-axis represents the possible return the agent can get for each episode, and the $y$-axis represents the percentage of getting a certain amount of return over $48,000$ episodes. The start state is uniformly distributed over all states except the goal. For each episode, the agent starts with a non-sink state, takes $20,000$ steps, and collects all rewards during the path. When $\alpha=0.6$, the return is in $(-1,2]$, and its mean value is $-0.074$, and its standard deviation is $0.228$. When $\alpha=0.2$, the return is in $(0,2]$, its mean value is $1.92$, and its standard deviation is $0.228$. Overall, Figure 4 shows that for the lower value of $\alpha$, the agent has a higher probability of avoiding falling off the cliff.
Third, to assess the stability of our EVaR-TRC Q-learning algorithm, we use six random seeds to generate samples, compute the optimal policies, and calculate the EVaR values on the CW domain. In Figure 4, the risk level $\alpha$ is $0.3$, and the $y$ axis represents the mean value and standard deviation of the EVaR values calculated in samples generated by six random seeds. Figure 4 shows that our Q-learning algorithm converges to similar solutions across different numbers of samples. The standard deviation is $0.021$. Note that our Q-learning algorithm could converge to a different optimal EVaR policy depending on the learning rate and the number of samples.
@@
Finally, we evaluate the convergence of Algorithm 2, which approximates the EVaR value calculated from the linear programming(LP) su2025risk on CW and GR domains. The $x$ axis represents the number of samples, and the $y$ axis represents the difference in EVaR values computed by LP and Q-learning algorithms. In Figure 6, the risk level $\alpha$ is $0.3$, and the EVaR value difference decreases to zero when the number of samples is around 20,000 on the CW domain. In Figure 6, the risk level $\alpha$ is $0.4$, and the EVaR value difference decreases to zero when the number of samples is around 30,000 on the GR domain. We illustrate the convergence result for some $\alpha$ values, but the conclusion applies to any $\alpha$ value. Overall, our Q-learning algorithm converges to the optimal value function.
@@ 6 Conclusion and Limitations
An important limitation of this work is that our algorithms focus on the tabular setting and do not analyze the impact of value function approximation. However, the work lays out solid foundations for building approximation scalable reinforcement learning algorithms. In particular, a strength of Q-learning is that it can be coupled with general value function approximation schemes, such as deep neural networks. Although extending convergence guarantees to such approximate settings is challenging, our convergence results show that our Q-learning algorithm is a sound foundation for scalable, practical algorithms. Future work should include rigorous numerical evaluation of deep learning implementation of the Q-learning algorithms.
Our theoretical analysis has two main limitation which may preclude its use in some practical settings. First, we must choose the limits $z_{\min},z_{\max}$. If these limits are set too small, the algorithms may fail to compute a good solution, and if they are too large, the algorithms may be excessively slow to detect the divergence of the TRC criterion for large values of $\beta$. The second limitation is the need to select the parameter $\beta_{0}$ that guarantees the existence of a $\delta$-optimal EVaR solution. A natural choice for $\beta_{0}$ was proposed in su2025risk , but it requires several runs of Q-learning just to establish an appropriate $\beta_{0}$. Future work needs to develop a better methodology for establishing these parameters. However, it is not clear that a proper model-free solution is possible without making additional assumptions on the structure of the MDP.
@@ Appendix A Proofs of Section 3
A.1 Proof of Theorem 3.1
@@
The exponential value function $w^{\star}(s)$ is the unique solution to the exponential Bellman equation [43, theorem 3.3]. Therefore, for a state $s\in\mathcal{S}$ and the optimal policy $\pi$, state value function $v^{\pi}(s)$, and the exponential state value function $w^{\pi}(s)$, $\beta\in\mathcal{B}$, we have that
@@
$v^{\star}(s)=\max_{a^{\prime}\in\mathcal{A}}q^{\star}(s,a,\beta)$ (16)
Then $q^{\star}(s,a,\beta)$ exists. The value $q^{\star}$ is unique directly from the uniqueness of $v^{\star}$.
A.2 Proof of Lemma 3.2
@@
$\ell_{\beta}(\tilde{x}-y)$ $=\beta^{-1}(e^{-\beta\cdot(\tilde{x}-y)}-1)+(\tilde{x}-y)$ (17)
@@
$y$ $=-\beta^{-1}\log(\mathbb{E}\bigl{[}e^{-\beta\cdot\tilde{x}}\bigr{]})$

### v2

@@ 3.2 EVaR Q-learning Algorithm
In this section, we adapt our ERM Q-learning algorithm to compute the optimal policy for the static EVaR-TRC objective in (3). As mentioned in Section˜2, EVaR is preferable to ERM because it is coherent and closely approximates VaR and CVaR. The construction of an optimal EVaR policy follows the standard methodology proposed in prior work su2025risk ; hau2023entropic . We only briefly summarize it in this paper.
@@
$h(\pi,\beta)$ $\;=\;\operatorname{ERM}_{\beta}^{\pi,\mu}\left[q^{\pi}(\tilde{s}_{0},\tilde{a}_{0},\beta)\right]+\beta^{-1}\log(\alpha)$ (17)
@@ 1 Construct $\mathcal{B}(\beta_{0},\delta)$ as described in (18) ;
2 Compute $(\pi^{\star}_{\beta},\,h^{\star}(\beta))$ by Algorithm˜1 for each $\beta\in\mathcal{B}(\beta_{0},\delta)$ where $h^{\star}(\beta)=\max_{\pi\in\Pi}h(\pi,\beta)$;
@@
Algorithm˜2 summarizes the procedure for computing a $\delta$-optimal EVaR policy. Note that the value of $h(\pi,\beta)$ may be -$\infty$, indicating either that the ERM-TRC objective is unbounded for $\beta$ or that the Q-learning algorithm diverged by random chance.
@@
Note that Algorithm˜2 assumes a small $\beta_{0}$ as its input. The derivation of $\beta_{0}$ is shown in A.5. It is unclear whether obtaining a prior bound on $\beta_{0}$ without knowing the model is possible. We, therefore, employ the heuristic outlined in Algorithm˜3 that estimates a lower bound $x_{\min}$ and an upper bound $x_{\max}$ on the random variable of returns. Then, we set $\beta_{0}:=\nicefrac{{8\delta}}{{(x_{\max}-x_{\min})^{2}}}.$
@@ 4 Convergence Analysis
This section presents our main convergence guarantees for the ERM-TRC Q-learning and EVaR-TRC Q-learning algorithms. The proofs for this section are deferred to Appendix˜B.
We require the following standard assumption to prove the convergence of the proposed Q-learning algorithms. The intuition of ˜4.1 is that each state-action pair must be visited infinitely often. Note that when an individual episode terminates upon reaching the sink state, the agent restarts a new episode from a random initial state.
@@
The input to Algorithm˜1 and Algorithm˜2 satisfies that
@@
The following theorem shows that the proposed ERM-TRC Q-learning algorithm enjoys convergence guarantees comparable to standard Q-learning. As Remark˜2 states, $z_{\min}$ and $z_{\max}$ are chosen to ensure that $q$ values are bounded.
@@
For $\beta\in\mathcal{B}$, assume that the sequence $\left(\tilde{\eta}_{i}\right)_{i=0}^{\infty}$ and $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})_{i=0}^{\infty}$ used in Algorithm˜1 satisfies ˜4.1 and step size condition
@@
where $i\in\{i\in\mathbb{N}\mid(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\}$, if $\tilde{z}_{i}\in[z_{\min},z_{\max}]$ almost surely, then the sequence $(\tilde{q}_{i})_{i=0}^{\infty}$ produced by Algorithm˜1 convergences almost surely to $q_{\infty}$ such that $q_{\infty}=\hat{B}_{\beta}q_{\infty}$.
The proof of Theorem˜4.2 follows an approach similar to that in the proofs of standard Q-learning bertsekas1996neuro with three main differences. First, the algorithm converges for an undiscounted MDP, and $\hat{B}_{\beta}$ is not a contraction captured by (bertsekas1996neuro, , assumption 4.4), which is restated as ˜B.2. Instead, we show that $\hat{B}_{\beta}$ satisfies monotonicity in Lemma˜4.3. Second, ˜B.2 leads to a convergence result somewhat weaker than the results for a contraction. Then, a separate boundedness condition (bertsekas1996neuro, , proposition 4.6) is imposed. Third, using the exponential loss function $\ell_{\beta}$ in (9) requires a more careful choice of step-size than the standard analysis.
@@
The following lemma shows that the loss function $\ell_{\beta}$ in (9) is strongly convex and has a Lipschitz-continuous gradient. These properties are instrumental in showing the uniqueness of our value functions and analyzing the convergence of our Q-learning algorithms by analyzing them as a form of stochastic gradient descent.
@@
Corollary˜4.5, which follows immediately from Theorem˜3.5, shows that Algorithm˜2 enjoys convergence guarantees and converges to a $\delta$-optimal EVaR policy.
@@
For $\alpha\in(0,1)$ and $\delta>0$, assume that the sequence $(\tilde{\eta}_{i})_{i=0}^{\infty}$ and $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})_{i=0}^{\infty}$ used in Algorithm˜1 satisfies ˜4.1 and step size condition
@@
where $i\in\{i\in\mathbb{N}\mid(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\}$, if $\tilde{z}_{i}\in[z_{\min},z_{\max}]$. Then Algorithm˜2 converges to the $\delta$-optimal stationary policy $\pi^{\star}$ for the EVaR-TRC objective in (3) almost surely.
The proof of Corollary˜4.5 follows a similar approach to the proofs of Theorem˜4.2. However, we also need to show that one can obtain an optimal ERM policy for an appropriately chosen $\beta$ that approximates an optimal EVaR policy arbitrarily closely.
@@ 5 Numerical Evaluation
In this section, we evaluate our algorithms on two tabular domains: cliff walking (CW) andrew2018reinforcement and gambler’s ruin (GR) su2025risk . The source code is available at https://github.com/suxh2019/ERM_EVaR_Q. First, we evaluate Algorithm˜2 on the CW domain. In this problem, an agent starts with a random state (cell in the grid world) that is uniformly distributed over all non-sink states and walks toward the goal state labeled by $g$ shown in Figure˜2. At each step, the agent takes one of four actions: up, down, left, or right. The action will be performed with a probability of $0.91$, and the other three actions will also be performed separately with a probability of $0.03$. If the agent hits the wall, it will stay in its place. The reward is zero in all transitions except for the states marked with $c$, $d$, and $g$. The agent receives $2$ in state $g$, $0.004$ in state $d$, 0.5,0.6,0.7,0.8-0.5,-0.6,-0.7,-0.8 and 0.9-0.9 in cliff region marked with $c$ from the left to the right. If an agent steps into the cliff region, it will immediately be returned to the state marked with $b$. Note that the agent has unlimited steps, and the future rewards are not discounted.
@@
We use Algorithm˜3 to compute $c$ and $d$ in Remark˜2 and estimate the bounds of $\tilde{z}_{i}(\beta)$ in Algorithm˜1. Figure˜2 and Figure˜2 show the optimal policies for EVaR with risk level $\alpha=0.2$ and $\alpha=0.6$ separately. Since the optimal policy is stationary, it can be analyzed visually. The arrow direction indicates the optimal action to take in that state. Note that the optimal actions for states $c$ and $g$ are the same and are omitted. In Figure˜2, the agent moves right to the last column of the grid world and then moves down to reach the goal. When the agent is close to the cliff region, it moves up to avoid risk. In Figure˜2, different optimal actions are highlighted in blue. As we can see, for different risk levels $\alpha$, when the agent is near the cliff region, it exhibits consistent behaviors to avoid falling off the cliff. For the remaining states, the agent behaves differently.
@@
Figure 4: Mean and standard deviation of EVaR values with $\alpha=0.2$ on CW domain
Second, to understand the impact of risk aversion on the structure of returns, we simulate the two optimal EVaR policies over $48,000$ episodes and display the distribution of returns in Figure˜4. The $x$-axis represents the possible return the agent can get for each episode, and the $y$-axis represents the probability of getting a certain amount of return. For each episode, the agent takes $20,000$ steps and collects all rewards during the path. When $\alpha=0.6$, the return is in (1,2](-1,2], and its mean value is 0.074-0.074, and its standard deviation is $0.228$. When $\alpha=0.2$, the return is in $(0,2]$, its mean value is $1.92$, and its standard deviation is $0.228$. Overall, Figure˜4 shows that for the lower value of $\alpha$, the agent has a higher probability of avoiding falling off the cliff.
Third, to assess the stability of Algorithm˜2, we use six random seeds to generate samples, compute the optimal policies, and calculate the EVaR values on the CW domain. In Figure˜4, the risk level $\alpha$ is $0.2$, the $x$ axis represents the number of samples, and the $y$ axis represents the mean value and standard deviation of the EVaR values. The standard deviation is $0.015$, and our Q-learning algorithm converges to similar solutions across different numbers of samples. Note that our Q-learning algorithm could converge to a different optimal EVaR policy depending on the learning rate and the number of samples.
@@
Finally, we evaluate the convergence of Algorithm˜2, which approximates the EVaR value calculated from the linear programming(LP) su2025risk on CW and GR domains. The $x$ axis represents the number of samples, the $y$ axis represents the difference in EVaR values computed by LP and Q-learning algorithms, and the risk level $\alpha$ is $0.2$. Figure˜6 and Figure˜6 show that the EVaR value difference decreases to zero when the number of samples is around 20,000 on the CW and GR domains. We illustrate the convergence results for $\alpha=0.2$, but the conclusion applies to any $\alpha$ value. Overall, our Q-learning algorithm converges to the optimal value function.
@@ 6 Conclusion and Limitations
An important limitation of this work is that our algorithms focus on the tabular setting and do not analyze the impact of value function approximation. However, the proposed algorithm is not limited to tabular representations. It is a general model-free gradient-based method that can be combined with any differentiable value function approximators. This work lays out solid foundations for building scalable approximation reinforcement learning algorithms. In particular, a strength of Q-learning is that it can be coupled with general value function approximation schemes, such as deep neural networks. Although extending convergence guarantees to such approximate settings is challenging, our convergence results show that our Q-learning algorithm is a sound foundation for scalable, practical algorithms.
Our theoretical analysis has two main limitations that may preclude its use in some practical settings. First, we must choose the limits $z_{\min},z_{\max}$. If these limits are set too small, the algorithms may fail to compute a good solution, and if they are too large, the algorithms may be excessively slow to detect the divergence of the TRC criterion for large values of $\beta$. It may be possible to use existing TD algorithms to improve how quickly we detect the divergence. The second limitation is the need to select the parameter $\beta_{0}$ that guarantees the existence of a $\delta$-optimal EVaR solution. A natural choice for $\beta_{0}$, proposed in su2025risk , requires several runs of Q-learning just to establish an appropriate $\beta_{0}$.
Acknowledgments
We thank the anonymous reviewers for their detailed reviews and thoughtful comments, which significantly improved the paper’s clarity. This work was supported, in part, by NSF grants 2144601 and 2218063.
@@ Appendix A Proofs of Section 3
A.1 Standard Results
For the model-based approach, Theorem˜A.1 shows that for an infinite horizon, the optimal exponential value function $w^{\infty,\star}$ is attained by a stationary deterministic policy and is a fixed point of the exponential Bellman operator. Following this theorem, we show that the ERM state-action value function can be computed as the fixed point of the Bellman operator shown in Theorem˜3.1 in the main paper.
Theorem A.1 [44], Theorem 3.3).
Whenever $w^{\infty,\star}>$-$\infty$ there exists $\pi^{\star}=(d^{\star})_{\infty}\in\Pi_{SD}$ such that
$w^{\infty,\star}\;=\;w^{\infty}(\pi^{\star})\;=\;L^{d^{\star}}w^{\infty,\star},$
and $w^{\infty,\star}$ is the unique value that satisfies this equation.
Lemma A.2 [18], Lemma A.7).
The function $\beta\mapsto\operatorname{ERM}_{\beta}\left[X\right]$ for any random variable $X\in\mathbb{X}$ and $\beta>0$ is continuous and non-increasing
Lemma A.3 [18], Lemma A.8).
Let $X\in\mathbb{X}$ be a bounded random variable such that $x_{min}<X<x_{max}$ a.s. Then, for any risk level $\beta>0$, $\operatorname{ERM}_{\beta}\left[\cdot\right]$ can be bounded as
$\mathbb{E}[X]-\frac{\beta(x_{max}-x_{min})^{2}}{8}\leq\operatorname{ERM}_{\beta}\left[X\right]\leq\mathbb{E}[X]$
A.2 Proof of Theorem 3.1
@@ A.1 Standard Results
The exponential value function $w^{\star}(s)$ is the unique solution to the exponential Bellman equation [44, theorem 3.3]. Therefore, for a state $s\in\mathcal{S}$ and the optimal policy $\pi$, state value function $v^{\pi}(s)$, and the exponential state value function $w^{\pi}(s)$, $\beta\in\mathcal{B}$, we have that
@@
$v^{\star}(s)=\max_{a\in\mathcal{A}}q^{\star}(s,a,\beta)$ (21)
Then $q^{\star}(s,a,\beta)$ exists. The value $q^{\star}(s,a,\beta)$ is unique directly from the uniqueness of $v^{\star}$.
A.3 Proof of Proposition 3.2
@@
$\frac{\partial\ell_{\beta}(\tilde{x}-y)}{\partial y}$ $=\frac{\partial\left(\beta^{-1}(e^{-\beta\cdot(\tilde{x}-y)}-1)+(\tilde{x}-y)\right)}{\partial y}$ (22)
@@
[…]